# Obliviate: Efficient Unmemorization for Protecting Intellectual Property in Large Language Models

### v1

@@ Abstract
Recent copyright agreements between AI companies and content creators have highlighted the need for precise control over language models’ ability to reproduce copyrighted content. While existing approaches rely on either complete concept removal through unlearning or simple output filtering, we propose Obliviate, a novel post-training technique that selectively prevents verbatim reproduction of specific text while preserving semantic understanding.
Obliviate operates by selecting tokens within memorized sequences and modifying the model’s probability distribution to prevent exact reproduction while maintaining contextual understanding. We evaluate Obliviate on multiple large language models (LLaMA-3.1 8B, LLaMA-3.1-instruct 8B, Qwen-2.5-7B, and Yi-1.5 6B) across both synthetic memorization tasks and organic copyright content. Our results demonstrate that Obliviate achieves orders of magnitude reduction, e.g., 100x, in verbatim memorization while maintaining model performance within 1% of baseline on standard benchmarks (HellaSwag, MMLU, TruthfulQA, and Winogrande). This makes Obliviate particularly suitable for practical deployment scenarios where companies need to efficiently address copyright concerns in pretrained models without compromising their general capabilities.
@@ 1 Introduction
The rise of large language models (LLMs) has sparked intense debate about intellectual property rights in AI-generated content. Central to this debate is the fundamental question: When does an AI system’s output constitute a derivative work versus an infringing reproduction? Recent legal frameworks and industry agreements have begun to converge on this distinction through specific policies targeting verbatim content reproduction(Bailey, ). For instance, publishers are discussing how large an excerpt would be acceptable to use(IngramSpark, ; Friedman, ). Similarly, major AI companies are starting to implement strict limits on exact text reproduction, with policies that will prevent generating copyrighted content and others that stop generating verbatim text after the generation exceeds a specific number of words.
Traditional approaches to protecting intellectual property in LLMs have primarily relied on post-processing filters, alignment or aggressive unlearning techniques that attempt to remove copyrighted content entirely from the model. However, emerging copyright discussions/agreements seems to focus more narrowly on protecting specific expressions rather than underlying ideas or concepts. This shift suggests a more targeted technical challenge: instead of preventing models from learning concepts altogether, we need only prevent them from reproducing specific sequences verbatim—a distinction that enables more efficient solutions.
In this work, we propose Obliviate, a novel technique that enables selective unmemorization of exact sequences in pretrained language models while maintaining their semantic knowledge of the text. Although recent work like the GoldFish loss (Hans et al., 2024) function has shown promise in preventing verbatim memorization during training, our approach uniquely addresses the critical challenge of modifying already trained models, where the vast majority of deployment scenarios lie.
Obliviate intuitively works by identifying specific tokens within the text that need to be unmemorized and applying targeted modifications. More concretely, we apply the KL (Kullback-Leibler) divergence loss to adjust the model’s output distribution, reducing the probability that the target token will reproduce ($\mathcal{L}_{\texttt{forget}}$) while maintaining fluency and coherence. Target tokens are selected using a sliding stride to ensure comprehensive coverage, and replacement candidates are carefully constrained to maintain the model’s utility.
For non-target tokens –those we do not wish to unmemorize– we apply KL divergence to encourage the top-k token distribution to remain consistent ($\mathcal{L}_{\texttt{maintain}}$), ensuring the model retains its overall performance. This dual strategy effectively balances the competing objectives of unmemorization and utility preservation.
We evaluate Obliviate on both synthetic datasets, where we explicitly train models to deeply memorize specific sequences, and on original organic copyrighted datasets, such as excerpts from Harry Potter, Moby Dick, and other well-known texts that appear in Web corpus. Our evaluation focuses on multiple large language models with different architectures and sizes, including LLaMA-3.1 8B, LLaMA-3.1-Instruct 8B, Qwen-2.5 7B and Yi-1.5 6B.
Our results demonstrate that Obliviate achieves strong unmemorization performance. For instance, it reduces the longest common subsequence between the generated data and target copyright content from an average of 270-380 words to less than 12 words for the synthetic dataset, while maintaining the model’s utility with a negligible drop and sometimes improvement within $1\%$ across baseline benchmarks, including MMLU, TruthfulQA, Hellaswag, and Winogrande. These findings highlight the effectiveness of Obliviate in selectively unmemorizing targeted content while preserving the overall performance of the models.
In short, our key contributions are as follows.
@@
  • We propose Obliviate, the first efficient post-training unmemorization technique that selectively targets verbatim text generation without compromising model utility.
  • We provide comprehensive empirical validation across diverse LLMs and benchmarks, demonstrating the strong performance of Obliviate.
@@ 2.1 Types of Model Memorization
Large Language Models (LLMs) exhibit two distinct types of training data memorization:
Semantic Memorization. The first type is semantic memorization, where models learn and internalize concepts from training data. This form of memorization is generally desirable as it enables models to understand and reason about learned concepts.
Verbatim Memorization. The second type is verbatim memorization, where models store and can reproduce exact sequences from their training data. This form of memorization is particularly problematic for copyright protection and is the primary focus of our work. We formally define verbatim memorization in terms of a model’s ability to complete sequences given partial contexts. Specifically, for a sequence $\mathbf{s}=(s_{1},...,s_{n})$, we say it is $(\ell,\beta)$-memorized if there exists a prefix length $\ell$ such that:
$P_{M}(s_{i+\ell},s_{i+\ell+1},\ldots,s_{n}|s_{i:i+\ell-1})>\beta$
where $i$ denotes the starting location of the prefix, formally defined as ${i\sim U[1,n-\ell]}$ with $U[1,n-\ell]$ denoting uniform sampling of prefix positions, and $\beta\in[0,1]$ is a threshold parameter.
Following Carlini et al. (2021), we simplify the memorization definition to consider greedy decoding: given a prefix $s_{i:i+\ell-1}$, we measure the longest sequence $n$ that model $M$ generates that matches the original text:
$n^{*}=\max{n:\forall j\in[i+\ell,i+n],s_{j}=\operatorname*{arg\,max}_{x\in\mathcal{V}}P_{M}(x|s_{i:j-1})}$
where $\operatorname*{arg\,max}_{x\in\mathcal{V}}$ represents greedy token selection from the vocabulary $\mathcal{V}$. Furthermore, to ensure we thoroughly evaluate memorization, we also consider using the default temperature during generation
@@ 2.2 Threat Model
We focus on copyright protection in non-adversarial settings, aiming to prevent models from unintentionally generating or being directly prompted to generate copyrighted content during normal usage. While a sufficiently motivated attacker might still be able to coerce the model into producing such content, our primary objective is to prevent routine copyright infringements.
@@ 3.1 Memorization in Language Models
Understanding and measuring memorization in language models has been an active area of research. Carlini et al. (2023); Zhang et al. (2023) demonstrated that large language models can memorize and reproduce substantial portions of their training data. Kandpal et al. (2022) and Lee et al. (2022) showed that memorization is particularly prevalent in commonly used web-scraped datasets, with exact duplicates in training data significantly increasing the likelihood of memorization. These findings raise significant concerns about privacy and copyright protection in deployed language models.
@@ 3.2 Machine Unlearning
Machine unlearning approaches focus on removing specific data points or concepts from trained models, targeting both semantic and verbatim memorization. These approaches can be categorized into exact and approximate unlearning methods. Exact unlearning methods provide theoretical guarantees for complete removal of target sequences. Bourtoule et al. (2020) proposed an efficient retraining approach using data sharding. However, subsequent work(Chen et al., 2021) demonstrated that such methods remain vulnerable to membership inference attacks (Shokri et al., 2017; Salem et al., 2019) when adversaries have access to both the original and unlearned models. Approximate unlearning methods trade theoretical guarantees for computational efficiency. Graves et al. (2021) introduced Amnesiac Unlearning, which tracks model updates and their corresponding training batches, enabling selective removal by ignoring specific updates. Various other approximate techniques have been proposed (Guo et al., 2023; Golatkar et al., 2020; Wu et al., 2022; Mehta et al., 2022; Eldan & Russinovich, 2023), each offering different trade-offs between unlearning effectiveness and computational cost.
While the recent GoldFish loss (Hans et al., 2024) similarly focuses on verbatim memorization, it operates during the training phase and cannot be directly applied to pretrained models. Our work adapts and extends these concepts for post-training modification, introducing novel techniques for maintaining model utility while targeting specific memorized sequences.
@@ 3.3 Copyright Protection
Recent work on copyright protection in language models has explored several approaches. Watermarking techniques (Li et al., 2023; Sablayrolles et al., 2020; Sander et al., 2024) enable detection of copyrighted content by embedding identifiable markers in training data—if a model generates text containing these watermarks, it indicates training on copyrighted content. However, these approaches focus on detection rather than prevention of copyright violations. In practice, companies primarily rely on filtering mechanisms to detect and block the generation of copyrighted content, such as Azure’s content filters, and alignment techniques that teach models to not generate copyrighted content. Obliviate improves upon these approaches by modifying the model’s parameters directly with a targeted loss function, making it inherently resistant to generating copyrighted content during normal usage without requiring additional computational overhead during inference from classifiers or blocklists.
@@
Figure 1: Visualization of Obliviate. Top of the figure shows an excerpt from ”Harry Potter and the Goblet of Fire,” where tokens designated for unmemorization are highlighted, using a stride of 10. Below, the three-stage process of distribution transformation are shown: (a) the initial prediction distribution($\mathbf{z}$), representing the model’s original probabilities for the dlgpointing token; (b) the target distribution after removing the target token, which serves as the post-forgetting target (${\mathbf{z}{\setminus x_{\texttt{target}}}}$); and (c) the final normalized distribution ($\text{softmax}\left({\mathbf{z}{\setminus x_{\texttt{target}}}}\right)$) used for computing the forgetting loss $\mathcal{L}_{\texttt{forget}}$.
@@ 4.1 Problem Formulation
Consider a pretrained language model $M$ with parameters $\theta$ that has potentially memorized sensitive sequences/text $T$ from its training data. Our goal is to modify $M$ such that it cannot reproduce these sequences verbatim while maintaining its ability to generate contextually appropriate text, i.e., without affecting $M$’s utility.
This approach differs from unlearning, where the model $M$ is expected to completely forget the concept of $T$ and not only avoid verbatim reconstruction of $T$. In fact, we aim for Obliviate to retain the concept of $T$, meaning the model still understands what ”Harry Potter” is and can answer questions about it, but without reproducing verbatim text from the Harry Potter books.
We believe this definition of unmemorization is crucial as it simplifies the unlearning problem and aligns with how most AI companies are addressing copyright issues. For example, Anthropic is discussing making deals with copyright owners to prevent exact reconstruction of their content(Brittain, ).
@@ 4.2 Obliviate
We take inspiration from the “GoldFish” loss (Hans et al., 2024) and adapt it to post-training, i.e., where the model is already trained and produces memorized –copyright-protected– content. Intuitively, Obliviate aims to perturb certain tokens of the memorized text $T$ to divert the model $M$ from outputting $T$, thereby generating entirely different text.
@@
To achieve this, Obliviate operates by defining a stride $s$, which it sweeps over the complete memorized text $T$ as a window. This stride $s$ is used to identify the target tokens for unmemorization, essentially skipping $s$ tokens after each target token to be unmemorized. For each of these selected tokens, we use the forget loss $\mathcal{L}_{\texttt{forget}}$ to forget these tokens, while employing the maintain loss $\mathcal{L}_{\texttt{maintain}}$ to preserve the behavior of the model. The top of Figure 1 shows an example of text from a Harry Potter book with the target tokens selected based on a stride of 10.
Forget Loss. $\mathcal{L}_{\texttt{forget}}$ employs KL-divergence (KL-div) on the output distribution of the target model, but with the top predicted token (the target token to unmemorize) removed. For this work, we select the top-10 probabilities to apply the KL-div and normalize them to compensate for the effect of removing the top token. The bottom of Figure 1 shows an example of the constructed distributions of a target token “ pointing”. More formally, $\mathcal{L}_{\texttt{forget}}$ is defined as follows:
$\mathcal{L}_{\texttt{forget}}=\text{KL}\left(\text{softmax}\left({\mathbf{z}}\right)\Bigg{|}\Bigg{|}\text{softmax}\left({\mathbf{z}{\setminus x_{\texttt{target}}}}\right)\right)$
where $\mathbf{z}$ represents the top-10 logits of the target model before removing the top predicted token, $\mathbf{z}{\setminus x_{\texttt{target}}}$ represents the logits after removing the top predicted token (the target token to unmemorize), and $x_{\texttt{target}}$ is the target token to unmemorize.
@@
Token Selection. To ensure that the model’s utility remains uncompromised, we selectively retain certain tokens even if they have been identified for unmemorization. Specifically, we exclude special tokens such as “start of text” and “end of text” to prevent damage to the model’s templates. Furthermore, we also exclude tokens with a 100% probability, as these tokens lack suitable alternative candidates. In other words, any replacement token at this location would be equally probable with a probability of 0%, rendering the candidate selection random and potentially pushing the model towards an undesired random direction, thus affecting its utility.
Maintain Loss. $\mathcal{L}_{\texttt{maintain}}$ is designed to ensure that the overall behavior of the model remains unchanged. This is typically achieved by minimizing the kl-divergence difference in output distributions between the original and the modified models across non-target tokens. We opt for KL-divergence loss over the standard cross-entropy loss for the maintain loss, as using cross-entropy loss could result in further memorization of the non-target tokens. Therefore, KL-divergence loss is employed to maintain the consistency of the model’s behavior without inducing additional memorization as a by-product.

### v2

@@ Abstract
Recent copyright agreements between AI companies and content creators underscore the need for fine-grained control over language models’ ability to reproduce copyrighted text. Existing defenses—ranging from aggressive unlearning to simplistic output filters—either sacrifice model utility or inadequately address verbatim leakage. We introduce Obliviate, a lightweight post-training method that surgically suppresses exact reproduction of specified sequences while preserving semantic understanding. Obliviate first identifies memorized passages and then, for each target token, minimally adjusts the model’s output distribution via a Kullback–Leibler divergence penalty to drive down the probability of exact reproduction. Simultaneously, we enforce a consistency loss on non-target tokens to retain the model’s fluency and task performance. We evaluate Obliviate on four popular 6–8B-parameter models (LLaMA-3.1, LLaMA-3.1-Instruct, Qwen-2.5, and Yi-1.5) using synthetic memorization benchmarks and organic copyrighted excerpts (e.g., Moby Dick, Frankenstein, Alice in Wonderland and Les Miserables). Across all settings, Obliviate reduces verbatim recall by two orders of magnitude (e.g., from hundreds of words to fewer than 12) while degrading downstream accuracy by at most 1% on HellaSwag, MMLU, TruthfulQA, and Winogrande. Furthermore, we benchmark Obliviate aganist different unlearning and copyright techniques using the MUSE and CoTaEval benchmarks. These results position Obliviate as a practical, high-fidelity solution for copyright compliance in deployed LLMs.
@@ 1 Introduction
The rise of large language models (LLMs) has intensified debates over intellectual property and security in AI-generated text. Federal inquiries by the U.S. Copyright Office and the European Union’s proposed AI Act [1, 2] highlight the need to distinguish lawful derivative use from infringing reproduction. At the heart of this discussion lies a crucial distinction between derivative works and infringing reproductions when models emit verbatim sequences from their training data [3]. Moreover, adversarial extraction attacks reveal that training data generation not only risks copyright violation but can also leak private or sensitive information [4].
Recent legal frameworks and industry accords have begun to address this gap by targeting verbatim replication policies [5]. For example, publishers are negotiating acceptable excerpt lengths [6, 7], while leading AI vendors impose strict caps on exact text regeneration, either by blocking copyrighted passages outright or halting generation once a specified word count is reached.
Historically, safeguarding intellectual property in LLMs has relied on post-processing filters, alignment procedures, or unlearning techniques designed to remove copyrighted/memorized text. In contrast, emerging agreements emphasize protecting specific expressions, e.g., exact generation, rather than abstract ideas. This shift reframes the technical challenge: rather than preventing models from internalizing concepts, we must disable them from generating copyrighted/protected token sequences.
1.1 Contribution
In this paper, we introduce Obliviate, the first –to the best of our knowledge– post-training unmemorization method that selectively suppresses exact token sequences in LLMs while preserving their utility. While recent work —such as the GoldFish loss—has demonstrated promise in suppressing memorization during training[8], Obliviate addresses the more common scenario of already-deployed models. This allows for a quick patching tool to address copyright issues—or memorization problems in general—after deployment.
Intuitively, Obliviate makes the LLM “forget” multiple tokens within a target span—such as copyrighted content—so that generation inevitably diverges and the probability of reproducing the original text approaches zero. To achieve this, Obliviate first selects the target tokens using a fixed-stride sliding window. A larger window size results in changing fewer tokens, thereby better preserving the model’s utility, while smaller sizes lead to forgetting more tokens, thus enhancing unmemorization performance. For this work, we find that a window size of four tokens achieves an optimal balance between these two objectives. Nonetheless, the window size is a hyperparameter that can be tuned for different applications.
After selecting the target tokens, we introduce a novel loss function called the “forget” loss ($\mathcal{L}_{\texttt{forget}}$). For each flagged token, we remove its highest-probability entry from the model’s output distribution and then minimize the Kullback–Leibler divergence between the original top-(k) distribution and the modified one ($\mathcal{L}_{\texttt{dist}}$). We selectively choose the candidate tokens for these target tokens to ensure that certain properties, such as capitalization, are preserved. An additional term is incorporated to penalize any residual probability mass on the target token, further driving its likelihood toward zero, namely the probability loss ($\mathcal{L}_{\texttt{prob}}$), i.e.,
$\mathcal{L}_{\texttt{forget}}=\mathcal{L}_{\texttt{dist}}+\mathcal{L}_{\texttt{prob}}$
To maintain the overall model behavior, we concurrently apply a “maintain” loss, $\mathcal{L}_{\text{maintain}}$, which is also based on KL divergence but focuses on non-target tokens. Optimizing the combined objective
$\mathcal{L}=\mathcal{L}_{\text{forget}}+\mathcal{L}_{\text{maintain}}$
enables selective suppression of verbatim replication while retaining fluency and factual knowledge.
We extensively evaluate Obliviate on both synthetic and organic datasets across multiple architectures and scales, including LLaMA-3.1 8B, LLaMA-3.1-Instruct 8B, Qwen-2.5 7B, and Yi-1.5 6B. Our evaluation demonstrates that Obliviate significantly reduces the generation probability of copyrighted content, with the longest common string length dropping from 270-380 to under 12 for syntactic data. We observe similarly strong performance on organic content, such as naturally occurring copyrighted texts (e.g., Frankenstein, Moby Dick, Alice in Wonderland, and Les Miserables). Furthermore, we assess the impact of Obliviate on the utility of the models across standard benchmarks (MMLU, TruthfulQA, HellaSwag, Winogrande). Our results show that Obliviate maintains the utility of the models, with no drop in most cases and a maximum drop of around 1%. Finally, we benchmark Obliviate using various metrics such as CoTaEval [9] and MUSE [10], which compares different techniques for enforcing copyright protection. The results shows that Obliviate achives the best utility-unmemorization trade off across all different techniques, e.g., Gradient Ascent (Grad Ascent)[11], Gradient Difference (Grad Diff)[12], KL Minimization (KL)[13], Preference Optimization (PO)[14], negative preference optimization (NPO) [15], and SimNPO [16].
In summary, our contributions are:
@@ 1.1 Contribution
  • We propose Obliviate, a novel post-training unmemorization framework that targets verbatim passages without harming LLM utility.
  • We demonstrate Obliviate’s efficacy and scalability through extensive experiments on multiple LLM families and both synthetic and organic datasets.
1.2 Organization
This paper is organized as follows: First, we present the background and preliminaries (Section 2), where we also introduce our threat model. Then, we present the related work (Section 3), methodology (Section 4), and evaluation (Section 5) sections. Finally, we discuss Obliviate’s limitations in Section 6 and conclude in Section 7.
@@ 2.1 Types of Model Memorization
LLMs exhibit two fundamentally different patterns of memorization during training:
Semantic Memorization. Semantic memorization occurs when a model memorizes abstract concepts, relationships, or factual knowledge from its training corpus. This behavior is generally beneficial: it allows the model to draw on learned representations to perform reasoning, answer questions, and transfer knowledge to novel contexts.
Verbatim Memorization. By contrast, verbatim memorization refers to the model’s ability to reproduce exact fragments of text seen during training. While semantic memorization leads to useful generalization, verbatim memorization raises serious concerns for privacy, copyright, and data leakage. It is precisely this capacity to regurgitate training sequences, rather than to generalize over them, that we address in this work.
We formalize verbatim memorization in terms of a model’s ability to complete text sequences given only partial context. Let $\mathbf{s}=(s_{1},\dots,s_{n})$ be any token sequence drawn from the training data. We say $\mathbf{s}$ is $(\ell,\beta)$–memorized if there exists a prefix length $\ell$ such that:
$P_{M}\bigl{(}s_{i+\ell},s_{i+\ell+1},\dots,s_{n}\mid s_{i:i+\ell-1}\bigr{)}\;>\;\beta,$
where the prefix start index $i$ is sampled uniformly from $\{1,\dots,n-\ell\}$ and $\beta\in[0,1]$ is a probability threshold.
Following the approach of [4], we further simplify this definition by focusing on greedy decoding. Given a prefix $s_{i:i+\ell-1}$, let the model generate tokens one at a time by selecting
$s_{j}^{\mathrm{gen}}=\arg\max_{x\in\mathcal{V}}P_{M}\bigl{(}x\mid s_{i:j-1}\bigr{)}\quad\text{for }j=i+\ell,i+\ell+1,\dots$
and let $n^{*}\;=\;\max\bigl{\{}\,n:s_{i:j}^{\mathrm{gen}}=s_{i:j}\ \forall\,j\leq i+n\,\bigr{\}}$ denote the length of exact reproduction. To capture generation under typical settings, we also evaluate $n^{*}$ when sampling with the model’s default temperature, rather than strictly greedy decoding.
@@ 2.2 Threat Model
Our goal is to safeguard against inadvertent disclosure of copyrighted material in typical, non-adversarial settings. Under this threat model, we seek to minimize the model’s tendency to reproduce protected text. Although a determined attacker might still be able to coerce the model into producing such content, our primary objective is to prevent routine copyright infringements.
@@ 3.1 Memorization in Language Models
Understanding and measuring memorization in language models has been an active area of research. [17, 18] demonstrated that large language models can memorize and reproduce substantial portions of their training data. [19] and [20] showed that memorization is particularly prevalent in commonly used web-scraped datasets, with exact duplicates in training data significantly increasing the likelihood of memorization. These findings raise significant concerns about privacy and copyright protection in deployed language models.
@@ 3.2 Machine Unlearning
Machine unlearning approaches focus on removing specific data points or concepts from trained models, targeting both semantic and verbatim memorization. These approaches can be categorized into exact and approximate unlearning methods. Exact unlearning methods provide theoretical guarantees for complete removal of target sequences. [21] proposed an efficient retraining approach using data sharding. However, subsequent work[22] demonstrated that such methods remain vulnerable to membership inference attacks [23, 24] when adversaries have access to both the original and unlearned models. Approximate unlearning methods trade theoretical guarantees for computational efficiency. [25] introduced Amnesiac Unlearning, which tracks model updates and their corresponding training batches, enabling selective removal by ignoring specific updates. Various other approximate techniques have been proposed [26, 27, 28, 29, 30], each offering different trade-offs between unlearning effectiveness and computational cost. For this work, we compare against multiple unlearning techniques that are part of the MUSE [10] benchmark and the copyright protection benchmark CoTaEval [9], namely Gradient Ascent (Grad Ascent) [11], Gradient Difference (Grad Diff) [12], KL Minimization (KL) [13], Preference Optimization (PO) [14], Negative Preference Optimization (NPO) [15], and SimNPO [16].
While the recent GoldFish loss [8] similarly focuses on verbatim memorization, it operates during the training phase and cannot be directly applied to pretrained models. Our work adapts and extends these concepts for post-training modification, introducing novel techniques for maintaining model utility while targeting specific memorized sequences.
@@ 3.3 Copyright Protection
Recent work on copyright protection in language models has explored several approaches. Watermarking techniques [31, 32, 33] enable detection of copyrighted content by embedding identifiable markers in training data—if a model generates text containing these watermarks, it indicates training on copyrighted content. However, these approaches focus on detection rather than prevention of copyright violations. In practice, companies primarily rely on filtering mechanisms to detect and block the generation of copyrighted content, such as Azure’s content filters, and alignment techniques that teach models to not generate copyrighted content. Obliviate improves upon these approaches by modifying the model’s parameters directly with a targeted loss function, making it inherently resistant to generating copyrighted content during normal usage without requiring additional computational overhead during inference from classifiers or blocklists.
@@
Figure 1: Visualization of Obliviate. Top of the figure shows an excerpt from "Harry Potter and the Goblet of Fire," where tokens designated for unmemorization are highlighted, using a stride of 10. Below, the three-stage process of distribution transformation are shown: (a) the initial prediction distribution($\mathbf{z}$), representing the model’s original probabilities for the dlgpointing token; (b) the target distribution after removing the target token, which serves as the post-forgetting target (${\mathbf{z}{\setminus x_{\texttt{target}}}}$); and (c) the final normalized distribution ($\text{softmax}\left({\mathbf{z}{\setminus x_{\texttt{target}}}}\right)$) used for computing the forgetting loss $\mathcal{L}_{\texttt{forget}}$.
@@ 4.1 Problem Formulation
Let $M$ be a pretrained language model with parameters $\theta$, which may have memorized sensitive or copyrighted sequences $T$ from its training data. Our objective is to modify $M$ so that it no longer reproduces these sequences verbatim, while preserving its ability to generate contextually appropriate text—i.e., without degrading its utility.
This goal differs from classical unlearning, where the model is expected to completely forget the concept of $T$ and not only avoid verbatim reconstruction. Instead, Obliviate retains the underlying concept (for example, understanding what “Harry Potter” is and answering related questions) but prevents exact reproduction of text from the Harry Potter books.
We believe that this notion of unmemorization simplifies the unlearning problem and aligns with industry practice. For instance, Anthropic has discussed agreements with copyright holders to block exact reconstruction of their content [34].
@@ 4.2 Obliviate
Building on the “GoldFish” loss [8], we adapt it to a post-training setting, where the model already exhibits memorized, copyrighted outputs. Intuitively, Obliviate perturbs selected tokens in the memorized text $T$ to divert $M$ away from generating $T$ exactly, causing it to produce alternative text.
@@
Stride and Token Selection. First, we define a stride $s$ that sweeps over the memorized text $T$ in fixed-size windows. This stride determines which tokens will be targeted for unmemorization essentially skipping $s$ tokens after each target token to be unmemorized. For each selected token, we apply a forget loss $\mathcal{L}_{\texttt{forget}}$ to reduce its likelihood, while concurrently using a maintain loss $\mathcal{L}_{\texttt{maintain}}$ to conserve the model’s overall behavior. The top of Figure 1 shows an example from a “Harry Potter” excerpt with tokens chosen using $s=10$.
@@
To ensure that the model’s utility remains uncompromised, we selectively retain certain tokens even if they have been identified for unmemorization. Specifically, we exclude special tokens such as “start of text” and “end of text” to prevent damage to the model’s templates. Furthermore, we also exclude tokens with a 100% probability, as these tokens lack suitable alternative candidates. In other words, any replacement token at this location would be equally probable with a probability of 0%, rendering the candidate selection random and potentially pushing the model towards an undesired random direction, thus affecting its utility.
4.2.1 Forget Loss
$\mathcal{L}_{\texttt{forget}}$ consists of two different components: the first ($\mathcal{L}_{\texttt{dist}}$) penalizes the output distribution of each target token, while the second ($\mathcal{L}_{\texttt{prob}}$) penalize the target tokens probability to further decrease them:
$\mathcal{L}_{\texttt{forget}}=\mathcal{L}_{\texttt{dist}}+\mathcal{L}_{\texttt{prob}}$
$\mathcal{L}_{\texttt{dist}}$ (Distribution Loss). Intuitively, the distribution loss use KL-divergence (KL-div) on the output distribution of the target model, but with the top predicted token (the target token to unmemorize) removed. For this work, we select the top-10 probabilities to apply the KL-div and normalize them to compensate for the effect of removing the top token. The bottom of Figure 1 shows an example of the constructed distributions of a target token “ pointing”. More formally, $\mathcal{L}_{\texttt{forget}}$ is defined as follows:
$\mathcal{L}_{\texttt{dist}}=\operatorname{KL}\Bigl{(}\operatorname{softmax}(\mathbf{z})\;\Big{\|}\;\operatorname{softmax}(\mathbf{z}\setminus x_{\mathrm{target}})\Bigr{)},$
where $\mathbf{z}$ represents the top-10 logits of the target model, $\mathbf{z}{\setminus x_{\texttt{target}}}$ represents the logits after removing the target token ($x_{\texttt{target}}$), and $x_{\texttt{target}}$ is the target token to unmemorize.
To select candidates for the $\mathcal{L}_{\texttt{forget}}$ loss, we inspect and filter the top-$k$ candidate tokens for each token targeted for unmemorization to ensure they are valid substitutions, i.e., to maintain the model’s utility, before applying the forget loss. As we iterate through the highest-probability tokens (excluding the original target token), we discard any candidates that meet one or more of the following criteria:
  1. 1.
    They are special tokens (e.g., end-of-text markers).
  2. 2.
    They violate formatting constraints: if the original token required a leading space or hyphen, the candidate must exhibit the same formatting.
  3. 3.
    They fail a capitalization check, ensuring that the replacement matches the original token’s casing pattern.
If filtering reduces the number of valid tokens below $k$, we extend our search to the next highest-probability tokens (i.e., from $k+1$, $k+2$, …) until we recover a full set of $k$ candidates for the forget loss.“‘
$\mathcal{L}_{\texttt{prob}}$ (Probability Loss). The second part of the forget loss, the probability loss ($\mathcal{L}_{\texttt{prob}}$), aims to further suppress the probability of each target token $p_{i}$. This results in the model being less probable to generate the target sequence/text. Concretely, the target loss ($\mathcal{L}_{\texttt{prob}}$) is defined as:
$\mathcal{L}_{\texttt{prob}}=\frac{1}{N}\sum_{i=1}^{N}\Bigl{[}\lambda_{1}\,\bigl{(}\log(1-p_{i}+\varepsilon)\bigr{)}^{2}\;+\;\lambda_{2}\,\log\!\bigl{(}\tfrac{p_{i}}{\tau}+\varepsilon\bigr{)}\,\tfrac{p_{i}}{\tau}\Bigr{]}$
where $\lambda_{1},\lambda_{2}$ weight the two terms, $\tau$ is a threshold ($10^{-4}$ in this paper), and $\varepsilon$ is a small constant for numerical stability. Intuitively, the first term pushes to reduce the probability to 0, while the second one add additional penalization to the probability if it is above specific threshold ($\tau$). The second term is set to zero when $p_{i}\leq\tau$.
Loss Scheduling. Naively combining the distribution ($\mathcal{L}_{\texttt{dist}}$) and probability ($\mathcal{L}_{\texttt{prob}}$) results in strong unmemorization of the target tokens, however as we saw in our experiments it degrades the utility of the models. After experimenting with different combination strategies, we found the best one is to first apply the distribution loss on its own until the model adequately forgets the target tokens. Then we combine both losses, i.e:
$\mathcal{L}_{\texttt{forget}}=\mathcal{L}_{\texttt{dist}}+\mathcal{L}_{\texttt{prob}}$
This way the model first forgets the target tokens while keeping its distribution mostly the same, then further push the target tokens in lower ranks without affecting the utility much (since $\mathcal{L}_{\texttt{dist}}$ already pushed the target tokens low enough and hence $\mathcal{L}_{\texttt{prob}}$ does not need to push for changing the model’s output distribution significantly). In short, this schedule ensures effective unmemorization while minimizing impact on the model’s utility.
4.2.2 Maintain Loss
The maintain loss $\mathcal{L}_{\texttt{maintain}}$ preserves the model’s overall behavior by minimizing the KL-divergence between the original and updated output distributions over non-target tokens. We favor KL-divergence over cross-entropy to avoid inadvertently reinforcing memorization of non-target tokens.